OpenStudy (anonymous):

An aircraft flies 400 km from a point O on a bearing of 025* and then 700 km on a bearing of 080* to arrive at B. a)how far north of O is B? b)how far east of O is B? c)find the distance and bearing of B from O?

OpenStudy (dumbcow):

 you can find distance x using law of cosines \[x^2 = 400^2 +700^2 - 2(400)(700) \cos 125\] you can find how far north using cosines \[\rightarrow 400 \cos 25 + 700 \cos 80\] how far east using sines \[\rightarrow 400 \sin 25 + 700 \sin 80\]

OpenStudy (perl):

i think you can use law of sines to find the bearing